OpenStudy (anonymous):

\[\left( -4a ^{5}b \right)^{2}\]

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\large \left( -4a ^{5}b \right)^{2}\] \[\large \left( -4\right)^2\left(a ^{5}\right)^2\left(b \right)^{2}\] ???

OpenStudy (anonymous):

-16(a5)2 (b)2

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Do you know the following rule:\[(ab)^2 = a^2b^2\] Do exactly that, think of -4a^5 as one term and b as the other. So using the rule gives us:\[(-4a^5b)^2=(-4a^5)^2b^2\] Now to get the (-4a^5)^2 part, we use the rule (ab)^2 = a^2b^2 once again. Now we treat -4 as one term and a^5 as the other term.\[(-4a^5b)^2=(-4a^5)^2b^2=(-4)^2(a^5)^2b^2=16a^{10}b^2\] @ilovedaivd

jimthompson5910 (jim_thompson5910):

(-4)^2 = (-4)*(-4) = 16 16 is positive, not negative

OpenStudy (anonymous):

wait